# Spin-dependent terms of the Breit-Pauli Hamiltonian evaluated with explicitly correlated Gaussian basis set for molecular computations

### v1

@@ IV Numerical tests of the implementation
This section presents numerical examples used to test the computer implementation of the spin-dependent Breit-Pauli Hamiltonian (BP) matrix elements. For the (atomic) test systems, relatively small basis sets are used, which can be readily extended for better-converged results. The value $\alpha^{-1}=137.$035 999 177, recommended by CODATA 2022 [57], is used in all computations.
@@ IV.1 Atomic test system
[…]
$\langle\hat{H}_{\text{Y}}\rangle_{0}$ $=2\langle\hat{H}_{\text{Y}}\rangle_{\pm 1}=-2\langle\hat{H}_{\text{Y}}\rangle_{\pm 2}$ (37)
$\langle\hat{H}_{\text{SS}}\rangle_{0}$ $=-2\langle\hat{H}_{\text{SS}}\rangle_{\pm 1}=10\langle\hat{H}_{\text{SS}}\rangle_{\pm 2},$ (38)
[…]
Table 1: Be $2\ ^{3}P^{\mathrm{o}}$ state: testing the implementation of the spin-dependent Breit–Pauli matrix elements. The non-relativistic energy, $E^{(0)}$ in $E_{\mathrm{h}}$, and the expectation value of spin-dependent operators for the $M_{\mathrm{J}}=0$ state in $\mu E_{\mathrm{h}}$, are given as a function of the number of ECG functions, $N_{\text{b}}$.
$N_{\text{b}}$ $E^{(0)}$ $\alpha^{2}\langle\hat{H}_{\text{SO}}\rangle_{0}$ $\alpha^{2}\langle\hat{H}_{\text{SOO}}+\hat{H}_{\text{SOO'}}\rangle_{0}$ $\alpha^{2}\langle\hat{H}_{\mathrm{SS,dp}}\rangle_{0}$
@@ IV.2 Molecular applications
Regarding molecular systems, we are not aware of any literature data for the high-precision computation of the spin-dependent BP matrix elements. Still, to be able to test the spin-dependent BP implementation reported in this work, we used the in-house implementation of the high-precision no-pair Dirac–Coulomb–Breit (DCB) approach [53, 54, 20, 21, 38, 58]. In particular, Ref. [38] is most relevant to this work. For the positive-energy projection, we employed the ‘cutting’ projector [54] and utilised quadruple precision arithmetic operations. The no-pair DCB approach is currently available only for two-electron systems; however, for two-electron triplet di- and triatomic systems, there is a non-vanishing contribution, which is used as a test case. The no-pair DCB energy includes not only the leading-order but also high-order $(Z\alpha)^{n}$ contributions; but, for small $Z$ nuclear charges, these higher-order contributions are small (so, we did not perform any ‘$\alpha$’ scaling of the variational result [54, 20, 59]).
@@
Table 2: $\mathrm{H}_{2}$${}^{3}\Sigma_{\text{g}}^{+}$ and b ${}^{3}\Sigma_{\text{u}}^{+}$ states ($R=1.4$ bohr): the non-relativistic energy, $E^{(0)}$ in $E_{\text{h}}$, and the zero-field splitting, $\Delta E$ in $\mu E_{\mathrm{h}}$. For the perturbative correction $\Delta E^{\mathrm{BP}}=\alpha^{2}\Delta E^{(2)}=\alpha^{2}[E^{(2)}_{0}-E^{(2)}_{\pm 1}]=\alpha^{2}[\langle H_{\mathrm{SS,dp}}\rangle_{0}-\langle H_{\mathrm{SS,dp}}\rangle_{\pm 1}]$ as a function of the number of fECGs, $N_{\text{b}}$. DCB superscript labels the no-pair DCB computations.
[…]
$N_{\text{b}}$ $E^{(0)}$ $\Delta E^{\mathrm{BP}}$ $\Delta E^{\mathrm{DCB}}$ $E^{(0)}$ $\Delta E^{\mathrm{BP}}$ $\Delta E^{\mathrm{DCB}}$
@@
Ref. [60] –0.713 0 0.02 –0.784 0 3.056
Triplet H+${}_{3}^{+}$
For the lowest-energy triplet state of $\mathrm{H}_{3}^{+}$, two nuclear geometries are considered (Table 3). Previous FCI computations [61, 62] converged the non-relativistic to 10–100 $\mu E\mathrm{h}$, but we are not aware of any spin-dependent BP computations and zero-field splitting values published in the literature.
We report computations using the spin-dependent BP implementation (presented in this work) and verify the results against our existing no-pair DCB approach at two distinct nuclear structures. First, we consider the equilibrium structure, which is linear and has two proton-proton distances equal to $R_{\text{pp},1}=R_{\text{pp},2}=2.454$ bohr [61]. Due to the triatomic, linear geometry, the threefold degeneracy of the energy is lifted into a non-degenerate and a doubly degenerate pair of states with $M_{S}=0$ and $M_{S}=\pm 1$ (where we chose the axis defined by the three nuclei for the quantization axis for the electron spin). Next, we repeated the computation for another, non-linear geometry (which is near a saddle point [62] on the potential energy surface) with two proton-proton distances, $R_{\text{pp},1}=1.939$ bohr, $R_{\text{pp},2}=5.961$ bohr, and one proton-proton-proton angle $\angle_{\text{ppp}}=64.4^{\text{o}}$. In this case, the three-fold degeneracy of the energy is completely lifted; so, we report two energy splittings for the three energy levels in Table 3. The spin-dependent BP implementation is in excellent agreement with the results of the no-pair DCB approach. The latter is currently applicable only for two-electron systems, so we proceed with the BP implementation for further molecular applications.
Table 3: Triplet $\mathrm{H}_{3}^{+}$ ground state: the non-relativistic energy, $E^{(0)}$ in $E_{\text{h}}$, and the zero-field splitting, $\Delta E$ in $\mu E_{\mathrm{h}}$, as a function of the number of fECGs, $N_{\text{b}}$ For linear (equilibrium) structure with centres of fECGs restricted to nuclear axis, $R_{\text{pp},1}=R_{\text{pp},2}=2.454$ bohr. For the general triangular structure the centers of the ECGs is restricted on the plane of the triangle and no projections are required, $R_{\text{pp},1}=1.939$ bohr, $R_{\text{pp},2}=5.961$ bohr, and $\angle_{\text{ppp}}=64.4^{\text{o}}$. For the equilibrium geometry, $\Delta E=E_{0}-E_{\pm 1}=\alpha^{2}[\langle\hat{H}_{\text{SS,dp}}\rangle_{0}-\langle\hat{H}_{\text{SS,dp}}\rangle_{\pm 1}]$. For the triangular structure, $\Delta E_{ij}=E_{i}-E_{j}$, where $i$ and $j$ labels the energy levels (please see also text). The DCB superscript labels the no-pair Dirac-Coulomb-Breit computation.
@@
$N_{\text{b}}$ $E^{(0)}$ $\Delta E^{\mathrm{BP}}$ $\Delta E^{\mathrm{DCB}}$ $E^{(0)}$ $\Delta E^{\mathrm{BP}}_{21}$ $\Delta E^{\text{DCB}}_{21}$ $\Delta E^{\mathrm{BP}}_{32}$ $\Delta E^{\text{DCB}}_{\mathrm{32}}$
@@ V Numerical applications: helium dimer
We compute the spin-dependent Breit-Pauli matrix elements connecting electronically excited states of the triplet He2 molecule (Fig. 1). The convergence of the non-relativistic energy at $\rho=2$ bohr is reiterated in Table 4, the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ state taken from Ref. [46] and the $\text{b}\ ^{3}\Pi_{\text{g}}$ and $\text{c}\ ^{3}\Sigma_{\text{g}}^{+}$ states taken from Ref. [47]. The $\text{B}\ ^{1}\Pi_{\text{g}}$ and $\text{C}\ ^{1}\Sigma_{\text{g}}^{+}$ states are computed in this work in order to account for the most important relativistic couplings of the a,b,c subspace. The electronic energy was converged by variational (non-linear) optimisation of the fECG basis functions using the stochastic variational method in combination with the Powell non-linear optimizer [9, 10, 63, 46].
In what follows, the atomic nuclei are along the (body-fixed) $z$ axis. The $z$ projection of the total, orbital plus electron spin, angular momentum, $\hat{J}_{z}=\hat{L}_{z}+\hat{S}_{z}$ is conserved for clamped nuclei, and its quantum number is labelled with $\Omega$ according to the spectroscopic practice [64]. Application of these results in rovibrational and rovibronic computations is reported in Ref. [47] and also in Ref. [46].
@@
Table 4: Triplet He2 at $\rho=2$ bohr: convergence of the the non-relativistic energy, $E_{x}^{(0)}$ in $E_{\text{h}}$, with respect to the number of optimised fECGs, $N_{x}$, for the $x=$ a, b, c, B, and C electronic states (Fig. 1). The ‘a’ state basis set is from Ref. [46], the ‘b’ and ‘c’ states are from Ref. [47].
@@
$N_{\mathrm{a}}$ $E_{\mathrm{a}}$ $N_{\mathrm{b}}$ $E_{\mathrm{b}}$ $N_{\mathrm{c}}$ $E_{\mathrm{c}}$ $N_{\mathrm{B}}$ $E_{\mathrm{B}}$ $N_{\mathrm{C}}$ $E_{\mathrm{C}}$
@@ V.1 Fine structure of the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ state
The lowest-energy triplet state of the $\mathrm{He}_{2}$ molecule is the a ${}^{3}\Sigma_{\text{u}}^{+}$ state. The relativistic magnetic dipole-dipole interaction, Eq. (11), lifts the threefold degeneracy of this state by splitting it into components with $\Omega=0$ and $\Omega=\pm 1$ (Fig. 2). This splitting can be characterized by a single parameter, $\kappa$, which is the energy deviation of the $\Omega=0$ state from the centroid. The corresponding shift of the $\Omega=\pm 1$ states is $-\frac{\kappa}{2}$, so the weighted average of the energies equals the centroid energy. The parameter $\kappa$ has relativistic and QED contributions,
@@
Since $\kappa^{(2)}$ originates solely from the relativistic magnetic dipole term (SS,dp), Eq. (11), its QED correction is obtained by simple multiplication, Eq. (15),
@@
[…]
The A ${}^{1}\Sigma_{\text{g}}^{+}$ state is closest to ‘a’ near equilibrium (Fig. 1), but there is no direct (first-order) coupling between a and A states through the relativistic (and QED) operators, Eqs. (16). Well, they can couple through second-order effects (some details are collected in the Supporting Information), exploratory computations show that these $\alpha^{4}E_{\text{h}}$ order contributions are tiny (on the few pico–Hartree, $\mathrm{p}E_{\mathrm{h}}$ level), so the A state is not considered further in this work.
@@ V.2 Fine structure of the $\text{b}\ ^{3}\Pi_{\text{g}}$ and $\text{c}\ ^{3}\Sigma_{\text{g}}^{+}$ states
Figure 2: Triplet $\mathrm{He}_{2}$: Splitting of the non-relativistic energies degenerate in $M_{S}$ by spin-dependent relativistic (and QED) terms. For the definition of $\lambda$, $\delta_{1}$, $\delta_{2}$, $\zeta,$ and $\xi$ please see Table 6 (and Ref. [47]). The colors indicate states, which couples with each other through relativistic interactions (red - $\lambda$, purple - $\delta_{2}$, green - $\delta_{1},\zeta,\xi$).
@@
Non-negligible couplings also exist with the singlet states, most importantly, with the $\text{B}\ ^{1}\Pi_{\text{g}}$ and the $\text{C}\ ^{1}\Sigma_{\text{g}}^{+}$ states (Fig. 1); the non-vanishing BP matrix elements are with $\lambda$, $\zeta$, and $\xi$ in Table 5. The computations of these matrix elements require all together about 500 CPU hours for the largest basis sets. Similarly to $\kappa$, the relativistic and QED couplings for the parameters $\beta$, $\gamma_{1}$, and $\epsilon$ differ only by a simple multiplicative factor, since only the magnetic dipole term contributes to the relativistic coupling,
@@
For the other matrix elements, the QED terms were calculated according to Eq. (15), and their convergence is shown in Table 7. For matrix elements with the singlet B and C states, due to the smaller basis sets used in this work, the non-relativistic energies are by 1-2 orders of magnitude less accurate than the a non-relativistic energy (Table 4). Despite this lower accuracy, the precision of the relativistic and QED couplings appears to be comparable to that of $\kappa$, the uncertainties are in the second digit after the decimal point (Tables 6 and 7). To test the convergence, we increased the B and C basis sizes to $N_{\text{B}}=500$ and $N_{\text{C}}=750$, respectively, while using the largest $b$ and $c$ basis sets. The resulting changes in the coupling values are in the 2nd to the 4th digits. These results are expected to be sufficiently accurate for the rovibronic computations, primarily focusing on b-c states, as presented in Ref. [47].
Table 5: Non-vanishing spin-dependent relativistic (and QED) matrix elements within the b, c, B, C electronic-spin subspace of He2. The superscript of each state labels the Cartesian component of the spatial part (as computed in QUANTEN) and the $M_{S}$ quantum number of the total electron-spin projection on the internuclear axis. The ‘.’ labels zero (0).
@@
$\text{b}^{x,-1}$ $-\beta$ $-\gamma_{1}$ $-\text{i}\gamma_{2}$ $-\text{i}\gamma_{1}$ $\delta_{2}$ $\lambda$
$\text{b}^{x,0}$ $2\beta$ $-\delta_{1}$ $\delta_{1}$ $\text{i}\zeta$
$\text{b}^{x,1}$ $-\gamma_{1}$ $-\beta$ $\text{i}\gamma_{1}$ $\text{i}\gamma_{2}$ $-\delta_{2}$ $\lambda$
$\text{b}^{y,-1}$ $\text{i}\gamma_{2}$ $-\text{i}\gamma_{1}$ $-\beta$ $\gamma_{1}$ $\text{i}\delta_{2}$ $\text{i}\lambda$
$\text{b}^{y,0}$ $2\beta$ $\text{i}\delta_{1}$ $\text{i}\delta_{1}$ $-\text{i}\zeta$
$\text{b}^{y,1}$ $\text{i}\gamma_{1}$ $-\text{i}\gamma_{2}$ $\gamma_{1}$ $-\beta$ $\text{i}\delta_{2}$ $-\text{i}\lambda$
$\text{c}^{0,-1}$ $-\delta_{1}$ $-\text{i}\delta_{1}$ $-\varepsilon$ $\xi$ $\text{i}\xi$
$\text{c}^{0,0}$ $\delta_{2}$ $-\delta_{2}$ $-\text{i}\delta_{2}$ $-\text{i}\delta_{2}$ $2\varepsilon$
$\text{c}^{0,1}$ $\delta_{1}$ $-\text{i}\delta_{1}$ $-\varepsilon$ $\xi$ $-\text{i}\xi$
$\text{B}^{x,0}$ $\text{i}\zeta$ $\xi$ $\xi$
$\text{B}^{y,0}$ $-\text{i}\zeta$ $-\text{i}\xi$ $\text{i}\xi$
$\text{C}^{0,0}$ $\lambda$ $\lambda$ $-\text{i}\lambda$ $\text{i}\lambda$
Table 6: Convergence of the relativistic spin-dependent contributions, in $\text{m}E_{\mathrm{h}}/\alpha^{2}$, within the a, b, c, B, and C electronic-spin subspace of He2 at $R=2.00$ bohr. $\kappa$ corresponds to the zero-field splitting of the ‘a’ state, and Table 5 defines all other non-vanishing matrix elements within the b, c, B, C subspace. $N_{\text{B}}=300$ and $N_{\text{C}}=500$, unless stated otherwise.
$N_{\mathrm{a}}$ $N_{\mathrm{b}}$ $N_{\mathrm{c}}$ $\kappa^{(2)}$ $\beta^{(2)}$ $\gamma^{(2)}_{1}$ $\gamma^{(2)}_{2}$ $\delta^{(2)}_{1}$ $\delta^{(2)}_{2}$ $\varepsilon^{(2)}$ $\lambda^{(2)}$ $\zeta^{(2)}$ $\xi^{(2)}$

### v2

@@ IV Numerical tests of the implementation
This section presents numerical examples used to test the computer implementation of the spin-dependent Breit-Pauli Hamiltonian (BP) matrix elements. For the (atomic) test systems, relatively small basis sets are used, which can be readily extended for better-converged results (if required). The value $\alpha^{-1}=137.$035 999 177, recommended by CODATA 2022 [69], is used in all computations.
@@ IV.1 Atomic test system
For the spherically symmetric (atomic) case, there are high-accuracy ECG implementations already available in the literature, including the analytic angular prefactor corresponding to higher angular momentum states [39, 40, 41]. As a first test case of our implementation, intended for the more general molecular problem, we computed the spin-dependent BP matrix elements of the 2 $^{3}P^{\mathrm{o}}$ state of the Be atom. We computed the expectation value of the spin-dependent BP operator terms, Eq. (14), labelled as $\langle\hat{H}_{\text{X}}\rangle_{J}\ (\text{X}=\text{SO, SOO \& SOO', SS,dp})$ with the $M_{J}=0$ wave function (Table 1). The relative error due to the finite basis size is approximately the same for all three terms; the deviations from the reference values are attributed to the incompleteness of our basis set. We also note that the expectation value for other $M_{J}=\pm 1,\pm 2$ values of the $2\ ^{3}P$ spin subspace of this atomic system can be calculated from the $\langle\hat{H}_{\text{Y}}\rangle_{0}$ matrix elements through known relations, e.g., in Ref. 40.
Table 1: Be $2\ ^{3}P^{\mathrm{o}}$ state: testing the implementation of the spin-dependent Breit–Pauli matrix elements. The non-relativistic energy, $E^{(0)}$ in $E_{\mathrm{h}}$, and the expectation value of spin-dependent operators for the $M_{J}=0$ state in $\mu E_{\mathrm{h}}$, are given as a function of the number of ECG functions, NN.
NN $E^{(0)}$ $\alpha^{2}\langle\hat{H}_{\text{SO}}\rangle_{0}$ $\alpha^{2}\langle\hat{H}_{\text{SOO}}+\hat{H}_{\text{SOO'}}\rangle_{0}$ $\alpha^{2}\langle\hat{H}_{\mathrm{SS,dp}}\rangle_{0}$
@@ IV.2 Molecular applications
Regarding molecular systems within the Born-Oppenheimer approximation, we are not aware of any literature data for the high-precision computation of the spin-dependent BP matrix elements. Still, to be able to test the spin-dependent BP implementation reported in this work, we used our in-house implementation of the high-precision no-pair Dirac–Coulomb–Breit (DCB) approach [66, 67, 20, 21, 48, 70]. In particular, Ref. 48 is most relevant to this work. For the positive-energy projection, we employed the ‘cutting’ projector [67] and utilised quadruple precision arithmetic operations. The no-pair DCB approach is currently available only for two-electron systems; however, for two-electron triplet di- and triatomic systems, there is a non-vanishing contribution, which is used as a test case. The no-pair DCB energy includes not only the leading-order but also high-order $(Z\alpha)^{n}$ contributions; but, for small $Z$ nuclear charges, these higher-order contributions are small (so, we did not perform any $\alpha$’ scaling of the variational result [67, 20, 71]).
@@
Table 2: $\mathrm{H}_{2}$$^{3}\Sigma_{\text{g}}^{+}$ and b $^{3}\Sigma_{\text{u}}^{+}$ states ($R=1.4$ bohr): the non-relativistic energy, $E^{(0)}$ in $E_{\text{h}}$, and the zero-field splitting, $\Delta E$ in $\mu E_{\mathrm{h}}$. For the perturbative correction $\Delta E^{\mathrm{BP}}=\alpha^{2}\Delta E^{(2)}=\alpha^{2}[E^{(2)}_{0}-E^{(2)}_{\pm 1}]=\alpha^{2}[\langle H_{\mathrm{SS,dp}}\rangle_{0}-\langle H_{\mathrm{SS,dp}}\rangle_{\pm 1}]$ as a function of the number of fECGs, NN. DCB superscript labels the no-pair DCB computations.
$^{3}\Sigma_{\text{g}}^{+} H2$$^{3}\Sigma_{\text{u}}^{+}$
NN $E^{(0)}$ $\Delta E^{\mathrm{BP}}$ $\Delta E^{\mathrm{DCB}}$ $E^{(0)}$ $\Delta E^{\mathrm{BP}}$ $\Delta E^{\mathrm{DCB}}$
@@
Ref. 72 –0.713 0 0.02 –0.784 0 3.056
[…]
For the lowest-energy triplet state of $\mathrm{H}_{3}^{+}$, two nuclear geometries are considered (Table 3). Previous FCI computations [73, 74] converged the non-relativistic energy to 10–100 $\mu E_{\mathrm{h}}$, but we are not aware of any spin-dependent BP computations and zero-field splitting values published in the literature.
We report computations using the spin-dependent BP implementation (presented in this work) and verify the results against our existing no-pair DCB approach at two distinct geometries. First, we consider the equilibrium structure, which is linear and has two proton-proton distances equal to $R_{\text{pp},1}=R_{\text{pp},2}=2.454$ bohr [73]. Due to the triatomic, linear geometry, the threefold degeneracy of the energy is lifted into a non-degenerate and a doubly degenerate pair of states with $M_{S}=0$ and $M_{S}=\pm 1$ (where we chose the axis defined by the three nuclei for the quantization axis of the electron spin). Next, we repeated the computation for another, non-linear geometry (which is near a saddle point [74] on the potential energy surface) with two proton-proton distances, $R_{\text{pp},1}=1.939$ bohr, $R_{\text{pp},2}=5.961$ bohr, and one proton-proton-proton angle $\angle_{\text{ppp}}=64.4^{\text{o}}$. In this case, the three-fold degeneracy of the energy is completely lifted; so, we report two energy splittings for the three energy levels in Table 3. The spin-dependent BP implementation is in excellent agreement with the results of the no-pair DCB approach. The latter is currently applicable only for two-electron systems, so we proceed with the BP implementation for further, poly-electronic molecular applications (with two or more clamped nuclei).
Table 3: Triplet $\mathrm{H}_{3}^{+}$ ground state: the non-relativistic energy, $E^{(0)}$ in $E_{\text{h}}$, and the zero-field splitting, $\Delta E$ in $\mu E_{\mathrm{h}}$, as a function of the number of fECGs, NN. For the linear (equilibrium) structure, the centres of the fECGs were restricted to the nuclear axis and $R_{\text{pp},1}=R_{\text{pp},2}=2.454$ bohr. For the general triangular structure, the centres of the fECGs were restricted in the plane of the triangle (without any further point-group symmetry projections), $R_{\text{pp},1}=1.939$ bohr, $R_{\text{pp},2}=5.961$ bohr, and $\angle_{\text{ppp}}=64.4^{\text{o}}$. For the linear (equilibrium) geometry, $\Delta E=E_{0}-E_{\pm 1}=\alpha^{2}[\langle\hat{H}_{\text{SS,dp}}\rangle_{0}-\langle\hat{H}_{\text{SS,dp}}\rangle_{\pm 1}]$. For the triangular structure, $\Delta E_{ij}=E_{i}-E_{j}$, where $i$ and $j$ labels the energy levels (please see also text). The DCB superscript labels the no-pair Dirac-Coulomb-Breit computation.
@@
NN $E^{(0)}$ $\Delta E^{\mathrm{BP}}$ $\Delta E^{\mathrm{DCB}}$ $E^{(0)}$ $\Delta E^{\mathrm{BP}}_{21}$ $\Delta E^{\text{DCB}}_{21}$ $\Delta E^{\mathrm{BP}}_{32}$ $\Delta E^{\text{DCB}}_{\mathrm{32}}$
@@ V Numerical applications: helium dimer
We compute the spin-dependent Breit-Pauli matrix elements connecting electronically excited states of the triplet He2 molecule (Fig. 1). The convergence of the non-relativistic energy at $\rho=2$ bohr (internuclear distance) is reiterated in Table 4; the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ state basis sets were taken from Ref. 26 and the $\text{b}\ ^{3}\Pi_{\text{g}}$ and $\text{c}\ ^{3}\Sigma_{\text{g}}^{+}$ states from Ref. 27. The $\text{B}\ ^{1}\Pi_{\text{g}}$ and $\text{C}\ ^{1}\Sigma_{\text{g}}^{+}$ states are computed in this work in order to account for the most important relativistic couplings of the a,b,c subspace. The electronic energy was converged by variational (non-linear) optimisation of the fECG basis functions using the stochastic variational method in combination with the Powell non-linear optimizer [9, 10, 64, 26].
In what follows, the atomic nuclei are along the (body-fixed) $z$ axis. The $z$ projection of the total, orbital plus electron spin, angular momentum, $\hat{L}_{z}+\hat{S}_{z}$ is conserved for clamped nuclei, and its quantum number is labelled with $\Omega$ according to the spectroscopic practice [75]. Application of these results in rovibrational and rovibronic computations is reported in Ref. 26 and also in Ref. 27.
@@
Table 4: He2 ($\rho=2$ bohr): the non-relativistic energy, $E_{x}^{(0)}$ in $E_{\text{h}}$, with respect to the number of optimised fECGs, $N_{x}$, for the $x=$ a, b, c, B, and C electronic states (Fig. 1). The ‘a’ state basis set is from Ref. 26, the ‘b’ and ‘c’ states are from Ref. 27.
@@
$N_{\mathrm{a}}$ $E^{(0)}_{\mathrm{a}}$ $N_{\mathrm{b}}$ $E^{(0)}_{\mathrm{b}}$ $N_{\mathrm{c}}$ $E^{(0)}_{\mathrm{c}}$ $N_{\mathrm{B}}$ $E^{(0)}_{\mathrm{B}}$ $N_{\mathrm{C}}$ $E_{\mathrm{C}}^{(0)}$
@@ V.1 Fine structure of the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ state
The lowest-energy triplet state of the $\mathrm{He}_{2}$ molecule is the a $^{3}\Sigma_{\text{u}}^{+}$ state. The relativistic magnetic spin dipole-dipole interaction, Eq. (11), lifts the threefold degeneracy of this state by splitting it into components with $\Omega=0$ and $\Omega=\pm 1$ (Fig. 2). This splitting can be characterized by a single parameter, $\kappa$, which is the energy deviation of the $\Omega=0$ state from the centroid. The corresponding shift of the $\Omega=\pm 1$ states is $\frac{\kappa}{2}$, so the weighted average of the energies equals the centroid energy. The parameter $\kappa$ has relativistic and QED contributions,
@@
Since $\kappa^{(2)}$ originates solely from the relativistic magnetic spin dipolar term (SS,dp), Eq. (11), its QED correction is obtained by simple multiplication, Eq. (15),
@@
For the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ state, we think that the non-relativistic energy is converged to 1–10 $\mu E_{\mathrm{h}}$ for the largest basis sets at $\rho=2$ bohr internuclear distance (Table 4). The convergence of $\kappa^{(2)}$ with respect to the number of basis functions is shown in Table 6. For the largest basis sets ($N_{\mathrm{a}}=1000$20002000), its value is converged to at least three significant digits. These results agree to one(-two) digit(s) of the Multireference Configuration Interaction (MRCI) computations reported in the literature [55]. The somewhat lower accuracy of the MRCI results is most probably related to the finite basis set error, which we can significantly improve by using an explicitly correlated basis set.
The A $^{1}\Sigma_{\text{g}}^{+}$ state is closest to the a-state near the equilibrium structure (Fig. 1), but there is no direct (first-order) coupling between a and A states through the relativistic (and QED) operators, Eqs. (16). Well, they can couple through second-order effects (some details are collected in the Supporting Information), exploratory computations show that these $\alpha^{4}E_{\text{h}}$ order contributions are tiny (on the few pico–Hartree, $\mathrm{p}E_{\mathrm{h}}$ level), so the A state is not considered further in this work.
@@ V.2 Fine structure of the $\text{b}\ ^{3}\Pi_{\text{g}}$ and $\text{c}\ ^{3}\Sigma_{\text{g}}^{+}$ states
Figure 2: Triplet $\mathrm{He}_{2}$: splitting of the non-relativistic energies degenerate in $M_{S}$ by spin-dependent relativistic (and QED) terms. For the definition of $\lambda$, $\delta_{1}$, $\delta_{2}$, $\zeta,$ and $\xi$ please see Table 5 (and Ref. 27). Colour-coding is employed to indicate states that are coupled via relativistic interactions between distinct non-degenerate subspaces (orange: $\lambda$, red: $\delta_{2}$, blue: $\delta_{1},\zeta,\xi$).
@@
Non-negligible couplings also exist with the singlet states, most importantly, with the $\text{B}\ ^{1}\Pi_{\text{g}}$ and the $\text{C}\ ^{1}\Sigma_{\text{g}}^{+}$ states (Fig. 1); the non-vanishing BP matrix elements are labelled with $\lambda$, $\zeta$, and $\xi$ in Table 5. Similarly to $\kappa$, the relativistic and QED couplings for the parameters $\beta$, $\gamma_{1}$, and $\epsilon$ differ only by a simple multiplicative factor, since only the magnetic spin dipolar term contributes to the relativistic coupling,
@@
For the other matrix elements, the QED terms were calculated according to Eq. (15), and their convergence is shown in Table 7. For matrix elements with the singlet B and C states, due to the smaller basis sets used in this work, the non-relativistic energies are by 1-2 orders of magnitude less accurate than the a non-relativistic energy (Table 4). Despite this lower accuracy, the precision of the relativistic and QED couplings appears to be comparable to that of $\kappa$, the uncertainties are in the second digit after the decimal point (Tables 6 and 7). To test the convergence, we increased the B and C basis sizes to $N_{\text{B}}=500$ and $N_{\text{C}}=750$, respectively, while using the largest b and c basis sets. The resulting changes in the coupling values are in the 2nd to the 4th digits. These results are expected to be sufficiently accurate for the rovibronic computations, primarily focusing on b-c states, as presented in Ref. 27.
Table 5: Non-vanishing spin-dependent relativistic (and QED) matrix elements within the b, c, B, C electronic-spin subspace of He2. The superscript of each state labels the Cartesian component of the spatial part (as computed in QUANTEN) and the $M_{S}$ quantum number of the total electron-spin projection on the internuclear axis. The ‘.’ labels zero (0). (Further details can be found in the Supplementary Material of Ref. 27.)
@@
$\text{b}^{x,-1}$ $\beta$ . $\gamma_{1}$ $\text{i}\gamma_{2}$ . $\text{i}\gamma_{1}$ . $\delta_{2}$ . . . $\lambda$
$\text{b}^{x,0}$ . $2\beta$ . . . . $\delta_{1}$ . $\delta_{1}$ . $\text{i}\zeta$ .
$\text{b}^{x,1}$ $\gamma_{1}$ . $\beta$ $\text{i}\gamma_{1}$ . $\text{i}\gamma_{2}$ . $\delta_{2}$ . . . $\lambda$
$\text{b}^{y,-1}$ $\text{i}\gamma_{2}$ . $\text{i}\gamma_{1}$ $\beta$ . $\gamma_{1}$ . $\text{i}\delta_{2}$ . . . $\text{i}\lambda$
$\text{b}^{y,0}$ . . . . $2\beta$ . $\text{i}\delta_{1}$ . $\text{i}\delta_{1}$ $\text{i}\zeta$ . .
$\text{b}^{y,1}$ $\text{i}\gamma_{1}$ . $\text{i}\gamma_{2}$ $\gamma_{1}$ . $\beta$ . $\text{i}\delta_{2}$ . . . $\text{i}\lambda$
$\text{c}^{0,-1}$ . $\delta_{1}$ . . $\text{i}\delta_{1}$ . $\varepsilon$ . . $\xi$ $\text{i}\xi$ .
$\text{c}^{0,0}$ $\delta_{2}$ . $\delta_{2}$ $\text{i}\delta_{2}$ . $\text{i}\delta_{2}$ . $2\varepsilon$ . . . .
$\text{c}^{0,1}$ . $\delta_{1}$ . . $\text{i}\delta_{1}$ . . . $\varepsilon$ $\xi$ $\text{i}\xi$ .
$\text{B}^{x,0}$ . . . . $\text{i}\zeta$ . $\xi$ . $\xi$ . . .
$\text{B}^{y,0}$ . $\text{i}\zeta$ . . . . $\text{i}\xi$ . $\text{i}\xi$ . . .
$\text{C}^{0,0}$ $\lambda$ . $\lambda$ $\text{i}\lambda$ . $\text{i}\lambda$ . . . . . .
[…]
$N_{\mathrm{a}}$ $N_{\mathrm{b}}$ $N_{\mathrm{c}}$ $\alpha^{2}\kappa^{(2)}$ $\alpha^{2}\beta^{(2)}$ $\alpha^{2}\gamma^{(2)}_{1}$ $\alpha^{2}\gamma^{(2)}_{2}$ $\alpha^{2}\delta^{(2)}_{1}$ $\alpha^{2}\delta^{(2)}_{2}$ $\alpha^{2}\varepsilon^{(2)}$ $\alpha^{2}\lambda^{(2)}$ $\alpha^{2}\zeta^{(2)}$ $\alpha^{2}\xi^{(2)}$